# Q-learning with Posterior Sampling

### v2

@@ Appendix D Concentration results
The result directly follows form Lemma 17. ∎
Corollary 4.
@@
Let $k_{i}$ denote the index of the episode when $(s,a)$ was visited for the $i_{\text{th}}$ time at step $h$. Set $x_{i}=\alpha_{n}^{i}(r_{k_{i},h}-R_{h}(s,a)+V(s_{i+1})-P_{s_{i},a_{i}}\cdot V)$ and consider filtration $\mathcal{F}_{i}$ as the $\sigma$-field generated by all random variables in the history set $\mathcal{H}_{k_{i},h}$. $r_{k_{i},h}-R_{h}(s,a)+V(s_{i+1})-P_{s_{i},a_{i}}\cdot(V)\leq H+1$. Using the definition of the learning rate (Lemma 16 (b)), we have $\sum_{i}^{n}x^{2}_{i}\leq H(H+1)^{2}/n$. We apply Azuma-Hoeffding inequality (see Lemma 14) combined with a union bound over all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and all possible values of $n\leq K$ to get the following with probability at least $1-\delta$,
@@
For $x_{k,h}=P_{s_{k,h},a_{k,h}}\cdot(V^{*}_{h+1}-V^{\pi_{k}}_{h+1})-(V^{*}_{h+1}(s_{k,h+1})-V^{\pi_{k}}_{h+1}(s_{k,h+1}))$ and filtration set $\mathcal{H}_{k,h}$ where $k$ is the episode index, $\{x_{k,h},\mathcal{H}_{k,h}\}$ forms a martingale difference sequence with $|x_{k,h}|\leq H$. We complete the proof using Lemma 14 and a union bound. ∎
Lemma 11.
@@
To bound $\sum_{k=1}^{K}(\mathbb{E}[Z_{k}]-Z_{k})$, we apply Bernstein inequality for martingale, Lemma 15, with $K=1$, $d=2SA\log(K)$ to get the required result.
Lemma 12.
@@
From Lemma 12 and (10), we get the result. ∎
Corollary 6.
@@
Using Lemma 11, we have with probability $1-\delta$,
@@ Appendix E Technical Preliminaries
See 2
@@
Lemma 14 (Corollary 2.1 in Wainwright [2019]).
@@
Lemma 15 (Lemma A8 in Cesa-Bianchi and Lugosi [2006]).
@@
Lemma 17 (Gaussian tail bound).
@@ Appendix F Sharper regret using Bernstein concentration
F.1 Proof of Theorem 3
The main mathematical reasoning closely follows that in the proof of Theorem 2 of Jin et al. [2018] with specific differences arising due to constant probability optimism and the definition of $\overline{V}_{k,h}$. For any $k,h,s,a$ with $n=N_{k,h}(s,a)$, we have $v_{b}(n,h,s,a)\leq 64\frac{H^{3}}{n+1}$. Therefore, Corollary 2 and Lemma 6 apply as they are. Hence, we get with probability at least $1-\delta$ for all $h$
@@ F.1 Proof of Theorem 3
Due to our observation that $v_{b}(n,h,s_{k,h},a_{k,h})\leq 64\frac{H^{3}}{n+1}$, Lemma 11 and Corollary 6 hold, therefore we get from (40),
@@
where the last inequality follows from Lemma 12. Since we have Lemma 22, we can follow the steps in (C.16) of Jin et al. [2018] to get
@@
Finally, combining with Corollary 2, we complete the proof.
@@ F.2 Supporting lemma
Corollary 7 (Corollary of Lemma 4).
@@
Lemma 18 (Based on Lemma C.7 in  Jin et al. [2018]).
Suppose (45) in Corollary 7 holds. For any $h\in[H]$, let $\beta_{k,h}\coloneqq\overline{V}_{k,h}(s_{k,h})-V^{*}_{h}(s_{k,h})$ and let $w=(w_{1},\ldots,w_{k})$ be non-negative weight vectors, then we have with probability at least $1-\delta$,
@@
(from Lemma 9 with probability at least $1-2\delta$)
@@
(from Corollary 7 with probability $1-\delta$)
@@
We now compute the summation $\sum_{k=1}^{K}w_{k}\beta_{k,h}$. We follow the proof of Lemma C.7 in Jin et al. [2018], with the only difference being the term $\sum_{k=1}^{K}\frac{w_{k}}{p_{2}}F(k,h,\delta)$, which we bound below. From the proof of Corollary 6.
@@
Lemma 19 (Based on Lemma C.3 of Jin et al. [2018]).
For any episode $k\in[K]$ with probability $1-\delta/K$, if Corollary 7 holds for all $k^{\prime}<k$, the for all $s,a,h\in\mathcal{S}\times\mathcal{A}\times[H]$:
@@
The proof is almost identical to that of Lemma C.3 of Jin et al. [2018] except we Lemma 18 instead of Lemma C.7 of Jin et al. [2018]. ∎
Lemma 20.
@@
Let $k_{i}$ denote the index of the episode when $(s,a)$ was visited for the $i_{\text{th}}$ time at step $h$. Set $x_{i}=\alpha_{n}^{i}(r_{k_{i},h}-R_{h}(s,a))$, $y_{i}=\alpha_{n}^{i}(V(s_{k_{i+1}})-P_{s_{i},a_{i}}\cdot V)$ and consider filtration $\mathcal{F}_{i}$ as the $\sigma$-field generated by all random variables in the history set $\mathcal{H}_{k_{i},h}$. We apply Azuma-Hoeffding (Lemma 14) to calculate $|\sum_{i=1}^{n}x_{i}|$ and Azuma-Bernstein for $|\sum_{i=1}^{n}y_{i}|$. Consider with probability $1-\delta$,
@@
Using Lemma 19 and a suitable union bound, we have with probability $1-\delta$,
@@
Lemma 21 (Based on Lemma 8).
@@
    (Posterior distribution mean) For any given $s,a$, with probability at least $\Phi(-1)-\frac{\delta}{H}-\delta$,
@@
    (Posterior distribution sample) For any given $s,a$, with probability at least $p_{1}$ ($p_{1}=\Phi(-1)-\frac{\delta}{H}-\delta$),
@@
    (In Algorithm 2) With probability at least $1-\frac{\delta}{KH}$,
@@
Here $\log(1/\delta^{\prime})=4$ and $\delta$ is a parameter of the algorithm used to define the number of samples $J$ used to compute the target $\overline{V}$.
@@
The proof is identical to that of Lemma 8 except that we use Lemma 20 instead of Corollary 4. ∎
Lemma 22 (Based on Lemma C.6 of Jin et al. [2018]).
@@
where for the first term we apply Lemma 19 (which holds when Corollary 7 and Lemma 18 hold) and for the second term we have from the definition of variance,

### v3

@@ Appendix D Concentration results
The result directly follows form Lemma E.6. ∎
Corollary D.2.
@@
Let $k_{i}$ denote the index of the episode when $(s,a)$ was visited for the $i_{\text{th}}$ time at step $h$. Set $x_{i}=\alpha_{n}^{i}(r_{k_{i},h}-R_{h}(s,a)+V(s_{i+1})-P_{s_{i},a_{i}}\cdot V)$ and consider filtration $\mathcal{F}_{i}$ as the $\sigma$-field generated by all random variables in the history set $\mathcal{H}_{k_{i},h}$. $r_{k_{i},h}-R_{h}(s,a)+V(s_{i+1})-P_{s_{i},a_{i}}\cdot(V)\leq H+1$. Using the definition of the learning rate (Lemma E.5 (b)), we have $\sum_{i}^{n}x^{2}_{i}\leq H(H+1)^{2}/n$. We apply Azuma-Hoeffding inequality (see Lemma E.3) combined with a union bound over all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and all possible values of $n\leq K$ to get the following with probability at least $1-\delta$,
@@
For $x_{k,h}=P_{s_{k,h},a_{k,h}}\cdot(V^{*}_{h+1}-V^{\pi_{k}}_{h+1})-(V^{*}_{h+1}(s_{k,h+1})-V^{\pi_{k}}_{h+1}(s_{k,h+1}))$ and filtration set $\mathcal{H}_{k,h}$ where $k$ is the episode index, $\{x_{k,h},\mathcal{H}_{k,h}\}$ forms a martingale difference sequence with $|x_{k,h}|\leq H$. We complete the proof using Lemma E.3 and a union bound. ∎
Lemma D.2.
@@
To bound $\sum_{k=1}^{K}(\mathbb{E}[Z_{k}]-Z_{k})$, we apply Bernstein inequality for martingale, Lemma E.4, with $K=1$, $d=2SA\log(K)$ to get the required result.
Lemma D.3.
@@
From Lemma D.3 and (10), we get the result. ∎
Corollary D.4.
@@
Using Lemma D.2, we have with probability $1-\delta$,
@@ Appendix E Technical Preliminaries
Lemma E.1 (High confidence from constant probability).
@@
Lemma E.3 (Corollary 2.1 in Wainwright [2019]).
@@
Lemma E.4 (Lemma A8 in Cesa-Bianchi and Lugosi [2006]).
@@
Lemma E.6 (Gaussian tail bound).
@@ Appendix F Sharper regret using Bernstein concentration
F.1 Proof of Theorem F.1
The main mathematical reasoning closely follows that in the proof of Theorem 2 of Jin et al. [2018] with specific differences arising due to constant probability optimism and the definition of $\overline{V}_{k,h}$. For any $k,h,s,a$ with $n=N_{k,h}(s,a)$, we have $v_{b}(n,h,s,a)\leq 64\frac{H^{3}}{n+1}$. Therefore, Corollary C.2 and Lemma 5 apply as they are. Hence, we get with probability at least $1-\delta$ for all $h$
@@
Due to our observation that $v_{b}(n,h,s_{k,h},a_{k,h})\leq 64\frac{H^{3}}{n+1}$, Lemma D.2 and Corollary D.4 hold, therefore we get from (41),
@@
where the last inequality follows from Lemma D.3. Since we have Lemma F.5, we can follow the steps in (C.16) of Jin et al. [2018] to get
@@
Finally, combining with Corollary C.2, we complete the proof.
@@ F.2 Supporting lemma
Corollary F.1 (Corollary of Lemma 4).
@@
Lemma F.1 (Based on Lemma C.7 in  Jin et al. [2018]).
Suppose (46) in Corollary F.1 holds. For any $h\in[H]$, let $\beta_{k,h}\coloneqq\overline{V}_{k,h}(s_{k,h})-V^{*}_{h}(s_{k,h})$ and let $w=(w_{1},\ldots,w_{k})$ be non-negative weight vectors, then we have with probability at least $1-\delta$,
@@
(from Lemma C.3 with probability at least $1-2\delta$)
@@
(from Corollary F.1 with probability $1-\delta$)
@@
We now compute the summation $\sum_{k=1}^{K}w_{k}\beta_{k,h}$. We follow the proof of Lemma C.7 in Jin et al. [2018], with the only difference being the term $\sum_{k=1}^{K}\frac{w_{k}}{p_{2}}F(k,h,\delta)$, which we bound below. From the proof of Corollary D.4.
@@
Lemma F.2 (Based on Lemma C.3 of Jin et al. [2018]).
For any episode $k\in[K]$ with probability $1-\delta/K$, if Corollary F.1 holds for all $k^{\prime}<k$, the for all $s,a,h\in\mathcal{S}\times\mathcal{A}\times[H]$:
@@
The proof is almost identical to that of Lemma C.3 of Jin et al. [2018] except we Lemma F.1 instead of Lemma C.7 of Jin et al. [2018]. ∎
Lemma F.3.
@@
Let $k_{i}$ denote the index of the episode when $(s,a)$ was visited for the $i_{\text{th}}$ time at step $h$. Set $x_{i}=\alpha_{n}^{i}(r_{k_{i},h}-R_{h}(s,a))$, $y_{i}=\alpha_{n}^{i}(V(s_{k_{i+1}})-P_{s_{i},a_{i}}\cdot V)$ and consider filtration $\mathcal{F}_{i}$ as the $\sigma$-field generated by all random variables in the history set $\mathcal{H}_{k_{i},h}$. We apply Azuma-Hoeffding (Lemma E.3) to calculate $|\sum_{i=1}^{n}x_{i}|$ and Azuma-Bernstein for $|\sum_{i=1}^{n}y_{i}|$. Consider with probability $1-\delta$,
@@
Using Lemma F.2 and a suitable union bound, we have with probability $1-\delta$,
@@
Lemma F.4 (Based on Lemma C.1).
@@
    (Posterior distribution mean) For any given $s,a$, with probability at least $1-\frac{2(k-1)\delta}{KH}-\frac{\delta}{KH}$,
@@
    (Posterior distribution sample) For any given $s,a$, with probability at least $p_{1}$ ($p_{1}=\Phi(-1)$),
@@
    (In Algorithm 2) With probability at least $1-\frac{2k\delta}{KH}$,
@@
The proof is identical to that of Lemma C.1 except that we use Lemma F.3 instead of Corollary D.2. ∎
Lemma F.5 (Based on Lemma C.6 of Jin et al. [2018]).
@@
where for the first term we apply Lemma F.2 (which holds when Corollary F.1 and Lemma F.1 hold) and for the second term we have from the definition of variance,
@@
Appendix G scratch
Define
$\sigma(N_{h}(s_{h},a_{h}))\coloneqq A_{h}(N_{h}(s_{h},a_{h}))+B_{h},$
where
$A_{h}(N_{h}(s_{h},a_{h}))=\sqrt{\frac{H^{3}\log(KH/\delta)}{N_{h}(s_{h},a_{h})}}$
and
$B_{h}\coloneqq\alpha_{N_{h}(s_{h},a_{h})}B_{h}+(1-\alpha_{N_{h}(s_{h},a_{h})})(\tilde{Q}_{h+1}(s_{h+1},a_{h+1})-\overline{V}_{h+1}(s_{h}+1)),$
which is same as,
$B_{k,h}=\sum_{i=1}^{N_{k,h}(s_{k,h},a_{k,h})}\alpha^{i}_{n}(\tilde{Q}_{k_{i},h+1}(s_{k_{i},h+1},a_{k_{i},h+1})-\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})).$